OpenStudy (anonymous):

using f(x)=x-2 and g(x)=5x+3 find f(g(x+2)) a. 5x+11 b. 5x+13 c.5x+5 d. -5x - 11 I think it's c, but i'm not really sure on how to solve this problem.

OpenStudy (anonymous):

easiest to work step by step

OpenStudy (anonymous):

g(x)=5x+3 g(x+2)=5(x+2)+3

OpenStudy (anonymous):

\[g(x)=5x+3\] \[g(x+2)=5((x+2)+3=5x+10+3=5x+13\]

OpenStudy (anonymous):

see how the change in f(x) changes the x in the function?

OpenStudy (anonymous):

I get it!

OpenStudy (anonymous):

It's like saying I replaced x with x+2 so I have to do the same in the function replace it.... for the second part you replace x with g(x+2) so replace x in f(x) with g(x+2) and everything else on the function as well.